# Comparison of pipeline, sequence-to-sequence, and GPT models for end-to-end relation extraction: experiments with the rare disease use-case

### v2

@@ 2.2 The three E2ERE methods
Generative pre-trained transformers (GPTs) have captured the fascination of the general public and researchers alike, especially since the introduction of ChatGPT in December 2022. However, the in-context learning and few-shot capabilities have already surfaced in June 2020, when Open AI released GPT-3 [25]. Building on the decoder component of the transformer architecture with the main objective of autoregressive left to right next token prediction task, they have excelled at text generation tasks (e.g., summarization). However, there is a growing interest in assessing their capabilities for language understanding tasks including relation extraction. BioGPT [26] and BioMedLM [27] have been pre-trained from scratch on biomedical abstracts from PubMed and full text articles from PubMed Central (from the corresponding subset from Pile [31]) based on the GPT-2 model [24]. In this effort, we focus on BioMedLM, a 2.7B parameter model, comprised of 32 layers, a hidden size of 2560, and 20 attention heads. BioMedLM is an order of magnitude larger than BioGPT and nearly twice as large as BioGPT. It has been shown to be be superior to BioGPT models (including in our experiments for this paper where BioGPT underperforms by 10-15% in F-score) and to our knowledge is the largest public GPT-style model for biomedicine. Hence, we only show BioMedLM results in this manuscript for the sake of clarify and simplicity. Unlike Seq2Rel whose sequence generation capabilities are highly constrained to terms observed in the input, BioMedLM and BioGPT are purely generative, and supervised fine-tuning involves using appropriate prompts and output templates. Technically, we could simply use the linearization schemas introduced for Seq2Rel. However, these generative models generate natural language statements and not unnatural-looking templates. So our initial experiments using a Seq2Rel style output schemas have failed. So, we considered two types of schemas here:
@@ 3 RESULTS AND DISCUSSION
Pipeline (SODNER+PURE): We used a batch size of 8, a learning rate of 1e-3, and 100 epochs to train the SODNER model for discontinuous entities with a PubMedBERT encoder. For the PURE NER model, we used PubMedBERT and trained for 100 epochs, with a learning rate of 1e-4 and a batch size of 8. We also experimented with PubMedBERT with the same settings. For the PURE relation model, we used both PubMedBERT and PubMedBERT as encoders with a learning rate of 1e-5 and trained for 25 epochs with the training batch size of 8.
Seq2Rel: Training was conducted for 150 epochs, with a learning rate of 2e-5 for the encoder (PubMedBERT or PubMedBERT) and 1.21e-4 for the decoder (LSTM) with a batch size of 2 and a beam size of 3 (for the decoder).
@@
The main results of the comparison using different models are presented in Table 4. For BioMedLM and T5 models, the ‘copyInstruct’ column in the table indicates the additional input prompt discussed earlier in this section where models are directed to only generate tokens observed in the input. We observe that the SODNER+PURE pipeline (with PubMedBERT encoder) produces the best F1-score of 52.2, which is 5 points more than the best-performing Seq2Rel model with the PubMedBERT encoder (47.15 F1), 5.2 points better than the best-performing model from T5 family (Flan-T5-large), and 13 points more than best performing BioMedLM model (38.9 F1). The pipeline’s performance does not increase when using the PubMedBERT model. For Seq2Rel, using PubMedBERT outperforms a model with PubMedBERT (44.53 F1) by 2.5 points, with an increase in both precision and recall. Potentially, the increased model capacity of PubMedBERT enables it to capture more complex and subtle relationships between medical terms and concepts. However, it is not clear why similar gains were not observed with PubMedBERT in the pipeline.
@@
SODNER + PURE | PubMedBERT | NA | 55.99 | 48.89 | 52.20 |
SODNER + PURE | PubMedBERT | NA | 56.20 | 48.52 | 52.08 |
Seq2Rel | PubMedBERT | NA | 47.60 | 40.90 | 44.53 |
PubMedBERT | NA | 51.46 | 43.51 | 47.15 | |
@@
Note that BioMedLM is over eight times larger than our best-performing pipeline model (considering it has three encoders based on the encoder PubMedBERT, which has 110M parameters). However, its low performance compared to the pipeline is not surprising because GPT models are autoregressive and do not benefit from language understanding arising from the bidirectional masked language modeling objective used in BERT models. Although the original BioMedLM [27] effort did not perform RE, it reports SOTA scores on biomedical Q&A tasks. The smaller BioGPT models were shown to do better than BERT models for E2ERE too. Hence we repurposed them for this RE task and as the largest publicly available GPT-based model, BioMedLM outperformed BioGPT models [26] by 10–15% in F1 score and we do not see these as worthy of reporting in this manuscript.

### v3

@@ 2.2 The three E2ERE methods
Generative pre-trained transformers (GPTs) have captured the fascination of the general public and researchers alike, especially since the introduction of ChatGPT in December 2022. However, the in-context learning and few-shot capabilities have already surfaced in June 2020, when Open AI released GPT-3 [25]. Building on the decoder component of the transformer architecture with the main objective of autoregressive left to right next token prediction task, they have excelled at text generation tasks (e.g., summarization). However, there is a growing interest in assessing their capabilities for language understanding tasks including relation extraction. BioGPT [26] and BioMedLM [27] have been pre-trained from scratch on biomedical abstracts from PubMed and full text articles from PubMed Central (from the corresponding subset from Pile [31]) based on the GPT-2 model [24]. In this effort, we focus on BioMedLM, a 2.7B parameter model, comprised of 32 layers, a hidden size of 2560, and 20 attention heads. BioMedLM is an order of magnitude larger than BioGPT and nearly twice as large as BioGPTlarge. It has been shown to be be superior to BioGPT models (including in our experiments for this paper where BioGPT underperforms by 10-15% in F-score) and to our knowledge is the largest public GPT-style model for biomedicine. Hence, we only show BioMedLM results in this manuscript for the sake of clarify and simplicity. Unlike Seq2Rel whose sequence generation capabilities are highly constrained to terms observed in the input, BioMedLM and BioGPT are purely generative, and supervised fine-tuning involves using appropriate prompts and output templates. Technically, we could simply use the linearization schemas introduced for Seq2Rel. However, these generative models generate natural language statements and not unnatural-looking templates. So our initial experiments using a Seq2Rel style output schemas have failed. So, we considered two types of schemas here:
@@ 3 RESULTS AND DISCUSSION
Pipeline (SODNER+PURE): We used a batch size of 8, a learning rate of 1e-3, and 100 epochs to train the SODNER model for discontinuous entities with a PubMedBERTbase encoder. For the PURE NER model, we used PubMedBERTbase and trained for 100 epochs, with a learning rate of 1e-4 and a batch size of 8. We also experimented with PubMedBERTlarge with the same settings. For the PURE relation model, we used both PubMedBERTbase and PubMedBERTlarge as encoders with a learning rate of 1e-5 and trained for 25 epochs with the training batch size of 8.
Seq2Rel: Training was conducted for 150 epochs, with a learning rate of 2e-5 for the encoder (PubMedBERTbase or PubMedBERTlarge) and 1.21e-4 for the decoder (LSTM) with a batch size of 2 and a beam size of 3 (for the decoder).
@@
The main results of the comparison using different models are presented in Table 4. For BioMedLM and T5 models, the ‘copyInstruct’ column in the table indicates the additional input prompt discussed earlier in this section where models are directed to only generate tokens observed in the input. We observe that the SODNER+PURE pipeline (with PubMedBERTbase encoder) produces the best F1-score of 52.2, which is 5 points more than the best-performing Seq2Rel model with the PubMedBERTlarge encoder (47.15 F1), 5.2 points better than the best-performing model from T5 family (Flan-T5-large), and 13 points more than best performing BioMedLM model (38.9 F1). The pipeline’s performance does not increase when using the PubMedBERTlarge model. For Seq2Rel, using PubMedBERTlarge outperforms a model with PubMedBERTbase (44.53 F1) by 2.5 points, with an increase in both precision and recall. Potentially, the increased model capacity of PubMedBERTlarge enables it to capture more complex and subtle relationships between medical terms and concepts. However, it is not clear why similar gains were not observed with PubMedBERTlarge in the pipeline.
@@
| SODNER + PURE | PubMedBERTbase | NA | 55.99 | 48.89 | 52.20 |
| SODNER + PURE | PubMedBERTlarge | NA | 56.20 | 48.52 | 52.08 |
| Seq2Rel | PubMedBERTbase | NA | 47.60 | 40.90 | 44.53 |
| PubMedBERTlarge | NA | 51.46 | 43.51 | 47.15 | |
@@
Note that BioMedLM is over eight times larger than our best-performing pipeline model (considering it has three encoders based on the encoder PubMedBERTbase, which has 110M parameters). However, its low performance compared to the pipeline is not surprising because GPT models are autoregressive and do not benefit from language understanding arising from the bidirectional masked language modeling objective used in BERT models. Although the original BioMedLM [27] effort did not perform RE, it reports SOTA scores on biomedical Q&A tasks. The smaller BioGPT models were shown to do better than BERT models for E2ERE too. Hence we repurposed them for this RE task and as the largest publicly available GPT-based model, BioMedLM outperformed BioGPT models [26] by 10–15% in F1 score and we do not see these as worthy of reporting in this manuscript.